i need help on this problem:
"the lengths of the sides of a triangle are three consecutive integers. writeand solve an equation to find the three integers. (hint:a^2+b^2=c^2)
How can i solve this problem??

Answer by mananth(16949)

"the lengths of the sides of a triangle are three consecutive integers. writeand solve an equation to find the three integers. (hint:a^2+b^2=c^2)

This must be a right triangle since a^2+b^2=c^2
..
let one leg be x
other leg = (x+1)
hypotenuse = (x+2)
(x+1)^2+x^2=(x+2)^2
x^2+2x+1+x^2=x^2+4x+4
x^2+2x+1-4x-4=0
x^2-2x-3=0
x^2-3x+x-3=0
x(x-3)+1(x-3)=0
(x-3)(x+1)=0
x= 3 OR x=-1
length cannot be negative
so x= 3
The sides are 3,4,5
...
m.ananth@hotmail.ca